# Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

### v1

@@ Abstract
Masked Image Modeling (MIM) has emerged as a pivotal approach for developing foundational visual models in the field of remote sensing (RS). However, current RS datasets are limited in volume and diversity, which significantly constrains the capacity of MIM methods to learn generalizable representations. In this study, we introduce RS-4M, a large-scale dataset designed to enable highly efficient MIM training on RS images. RS-4M comprises 4 million optical images encompassing abundant and fine-grained RS visual tasks, including object-level detection and pixel-level segmentation. Compared to natural images, RS images often contain massive redundant background pixels, which limits the training efficiency of the conventional MIM models. To address this, we propose an efficient MIM method, termed SelectiveMAE, which dynamically encodes and reconstructs a subset of patch tokens selected based on their semantic richness. SelectiveMAE roots in a progressive semantic token selection module, which evolves from reconstructing semantically analogical tokens to encoding complementary semantic dependencies. This approach transforms conventional MIM training into a progressive feature learning process, enabling SelectiveMAE to efficiently learn robust representations of RS images. Extensive experiments show that SelectiveMAE significantly boosts training efficiency by 2.2-2.7 times and enhances the classification, detection, and segmentation performance of the baseline MIM model. The dataset, source code, and trained models will be released at RS-4M.
@@ 1 Introduction
Over the past decade, advancements in remote sensing (RS) technology and data acquisition have significantly improved applications in ecosystem monitoring [1], natural disaster management [2], among others [3, 4]. These applications rely on essential capabilities such as scene classification [5, 6], object detection [7], change detection [8], and semantic segmentation [9]. However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
@@
Thanks to significant advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [10, 11], the pre-training of visual foundation models has seen remarkable improvements [12, 13, 14, 15, 16, 17]. Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [18]. However, two challenges persist in the development of RSFMs. (i) Compared to the ImageNet-21k [19] dataset, previous RS datasets [20, 21, 22, 23] contain significantly fewer samples (approximately 1 million vs. 14 million), which impedes the sufficient MIM training of large backbones. (ii) These datasets, which primarily focus on global scene semantics [20, 21, 22], lack the diversity and fine-grained information of RS scenarios encountered in downstream tasks. This limitation restricts the generalization of the learned representations.
@@
To address these challenges, we introduce a large-scale RS dataset called RS-4M. This dataset, comprising 4 million optical images, is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. RS-4M significantly exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger. Moreover, RS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 1 left).
@@
Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the significant computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [39] backbone on 8 Nvidia A100 GPUs [18]. In natural scene analysis, this issue has led to numerous studies [40, 41, 42, 43, 44, 45] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [40, 41]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [42, 43, 44], significantly speeding up feature extraction.
@@
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [18, 28]. This raises two key questions about how to efficiently conduct MIM training in the RS field: (i) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? (ii) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
@@
We propose an efficient MIM method called SelectiveMAE, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens ($e.g.,\leq$50%) with higher HOG values for feature encoding ($e.g.,\leq$15%) and pixel reconstruction ($e.g.,\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing (Fig. 3). To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically leverages semantically relevant patch tokens throughout the entire MIM training phase. At the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary approach allows SelectiveMAE to efficiently and progressively learn robust representations of RS images, thereby accelerating MIM convergence. Our experiments reveal that only 40% of RS image patches are sufficient to train a comparable MIM model, offering new insights into conducting MIM training on RS images.
The contributions of this paper are summarized as follows. (1) We introduce the RS-4M dataset, a large-scale optical remote sensing dataset for unsupervised learning methods to date, distinguished by its diverse scene details. (2) We propose SelectiveMAE, a novel and efficient MIM method tailored for remote sensing images. This method incorporates a new PSTS module, which significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
Experiment results suggest that compared to the baseline method, SelectiveMAE not only achieves 2.22.7×2.2-2.7$\times$2.2 - 2.7 × speedup in pre-training but also provides performance improvement of 5.6% mAP and 1.4% mF1 when applied to downstream object detection and segmentation tasks, respectively.
@@ 2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [23, 46], object detection [47, 48, 49], and segmentation [50, 51, 52]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works, like SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches, combine various sensor data to create extensive datasets, while many other datasets focus solely on optical RS images. MillionAID [20] offers a million-scale dataset for RS optical image classification, while SeCo [22] and CACo [21] provide nearly a million images of the same location over different times. These datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, we introduce the RS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
Remote Sensing Foundation Models. remains unlabeled and thus inaccessible for supervised learning [54]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Although there are some contrastive self-supervised methods [24, 21, 22, 55], recent advancements have primarily centered around generative self-supervised methods. For example, RVSA [18] leverages the vision transformer backbone pre-trained by the vanilla MAE [10] method and introduces computationally efficient rotated varied-size window attention to replace the original attention during fine-tuning. In addition, many studies are focusing on improving generative self-supervised algorithms by leveraging general image knowledge [25, 26, 28], scaling up parameter sizes [29], integrating spatio-temporal data [30, 31, 32], handling multi-sensor data [56, 57, 58, 59, 60], and employing multi-scale concepts [33, 35, 36]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In this paper, we propose a novel acceleration method that significantly speeds up training, enhancing practicality in the RS domain.
@@
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [61], MIM has been developed for visual pre-training [62, 11, 10, 63, 64]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [65, 66, 67, 68, 69], masking strategies [70, 71], and reconstruction methods [72, 73, 74, 75]. For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or ues a novel Difficulty-Flatten Loss [43]. Additionally, CrossMAE [45] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. In this paper, we introduce a novel PSTS module to address this issue and accelerate the training process.
3 Dataset
3.1 Data collection and organization
@@
Figure 1: Left: The organization of data sources in RS-4M. Right: The comparison between RS-4M and other datasets.
Recent advancements in self-supervised pre-training RSFMs are hindered by the smaller scale and diversity of existing RS datasets compared to natural scene datasets. To address this issue, we propose the RS-4M dataset, a continuously updated large-scale RS dataset following the principles of Diversity, Richness, and Scalability (DiRS) [20]. (1) We collected and organized datasets from the past decade, focusing on mainstream RS image interpretation tasks, including scene classification, object detection, semantic segmentation, and change detection. These datasets, detailed in Fig. 1, encompass around 100 sub-datasets with a total of 13.5 million images. (2) We excluded multispectral and SAR data, keeping only optical images in this study. Future updates will incorporate multispectral and SAR data to support multi-modal self-supervised pre-training. (3) Unlike datasets like MillionAID [20] designed for scene classification, our RS dataset is designed to facilitate downstream detection and segmentation tasks. To address large-scale image in segmentation datasets, we randomly cropped high-resolution images into smaller slices. (4) Images with pixels below 64 or above 1024 were removed to facilitate self-supervised pre-training. (5) We combined the remaining images and eliminated duplicates using a two-phase process: a coarse phase with perceptual hashing[76] and a refined phase with manual review. With the Hash value to select and the checking of manual review, only highly similar images are excluded. This resulted in the RS-4M dataset containing about 4 million high-quality RS optical images, which is four times larger than previous representative RS datasets, as shown in Fig. 1 right.
Table 1: Comparison between RS-4M and MillionAID: ViT Base models [39] are pre-trained on these datasets using MAE [10] and fine-tuned on the RESISC-45 dataset [77].
Dataset Images Number Epoch Top-1 Accuracy
MillionAID [20] 1 Million 800 89.20
RS-4M 2 Million 400 91.80
RS-4M 3 Million 267 92.24
RS-4M 4 Million 200 92.38
3.2 Preliminary Evaluation of RS-4M Dataset
RS-4M offers a significantly larger and more diverse image set compared to previous datasets. To evaluate its effectiveness, we pre-train a ViT base model [39] using the vanilla MAE method [10]. For comparison, we use the MillionAID dataset, maintaining an equal number of data points during training: 800 epochs for MillionAID’s 1 million images and 200 epochs for our RS-4M dataset. In addition, we create new datasets by sampling different numbers of images from RS-4M for further comparison. After pre-training, we fine-tune the models on the downstream RESISC-45 dataset [77] and assess the Top-1 classification accuracy. The results in Table 1 demonstrate that the RS-4M dataset outperforms MillionAID across various sample sizes, with the highest accuracy of 92.38% achieved using the entire dataset. Even with fewer images, RS-4M consistently delivers superior performance, highlighting its greater diversity and effectiveness. These findings underscore the advantage of using larger and more varied datasets for unsupervised pre-training.
4 Method
4.1 Masked Autoencoders Preliminaries

### v3

# OpticalRS-4M: Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

@@ Abstract
Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS). However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations. Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead. To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We curated a high-quality dataset named OpticalRS-4M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication. OpticalRS-4M comprises 4 million optical images covering various RS tasks, such as object detection and pixel segmentation. To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images. Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times. This highlights the effectiveness and scalability of our pipeline in developing RS foundational models. The dataset, source code, and trained models will be released at OpticalRS-4M.
@@ 1 Introduction
Over the past decade, advancements in remote sensing (RS) technology and data acquisition have improved applications in ecosystem monitoring [1], natural disaster management [2], among others [3, 4]. These applications rely on essential capabilities such as scene classification [5, 6], object detection [7], change detection [8], and semantic segmentation [9]. However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
@@
Thanks to advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [10, 11], the pre-training of visual foundation models has seen remarkable improvements [12, 13, 14, 15, 16, 17]. Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [18]. However, two challenges persist in the development of RSFMs. (i) Compared to the ImageNet-21k [19] dataset, previous RS datasets [20, 21, 22, 23] contain fewer samples (approximately 1 million vs. 14 million), which impedes the sufficient MIM training of large backbones. (ii) These datasets, which primarily focus on global scene semantics [20, 21, 22], lack the diversity and fine-grained information of RS scenarios encountered in downstream tasks. This limitation restricts the generalization of the learned representations.
@@
To address these challenges, we propose a new pipeline to collect, create, and efficiently process a large optical RS dataset. Firstly, we reviewed publicly available remote sensing datasets from the past decade and selected them based on the DiRS (Diversity, Richness, and Scalability) principle [20]. Nevertheless, issues such as inconsistent data sources, excessively large image sizes, and redundant pixels still exist. Therefore, we also apply exclusion, slicing, and deduplication processes to further improve data quality. Correspondingly, we obtain a large-scale RS dataset called OpticalRS-4M comprising 4 million optical images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. OpticalRS-4M exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger. Moreover, OpticalRS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 1 left).
@@
Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [39] backbone on 8 Nvidia A100 GPUs [18]. This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-4M. In natural scene analysis, this issue has led to numerous studies [40, 41, 42, 43, 44, 45] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [40, 41]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [42, 43, 44], speeding up feature extraction.
@@
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [18, 28]. This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? 2) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
@@
Regarding the above issues, in the second part of the pipeline, we introduce a MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens ($e.g.,\leq$50%) with higher HOG values for feature encoding ($e.g.,\leq$15%) and pixel reconstruction ($e.g.,\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing. To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase. In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence. Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
The contributions of this paper are summarized as follows:
1) We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models.
2) Using this pipeline, we create the OpticalRS-4M dataset, which is a large-scale RS dataset comprising 4 million optical images with diversified coverage scenarios.
3) We introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training. It significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
Experiment results suggest the effectiveness and scalability of the proposed pipeline. OpticalRS-4M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [10].
@@ 2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [23, 46], object detection [47, 48, 49], and segmentation [50, 51, 52]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works combine various sensor data to create extensive datasets. SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches. SSL4EO-S12 [54] dataset uses publicly available SAR/optical satellite data collected by the European Space Agency’s Sentinel mission. SSL4EO-L [55] consists entirely of multispectral data. SatlasPretrain [56] is constructed from data sources such as Sentinel-2, Sentinel-1, Landsat 8/9, and 0.5 - 2 meter/pixel aerial imagery. While it includes millions of images, the majority are multispectral and SAR, with only a small fraction being RGB. There are also some large-scale RS visual-language datasets such as Skyscript [57] and RS5M [58], but they primarily focus on multimodal tasks. Currently, some datasets focus solely on optical RS images. MillionAID [20] offers a million-scale dataset for RS optical image classification, while SeCo [22] and CACo [21] provide nearly a million images of the same location over different times. However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, with the proposed pipeline, we introduce the OpticalRS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
@@
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [59], MIM has been developed for visual pre-training [60, 11, 10, 61, 62]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [63, 64, 65, 66, 67], masking strategies [68, 69], and reconstruction methods [70, 71, 72, 73]. For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or use a novel Difficulty-Flatten Loss [43]. Additionally, CrossMAE [45] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. Considering these issues, we introduce an efficient MIM method named SelectiveMAE that significantly speeds up pre-training, enhancing the practicality of developing RS foundation models based on large-scale datasets.
Remote Sensing Foundation Models. Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [74]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Although there are some contrastive self-supervised methods [24, 21, 22, 75], recent advancements have primarily centered around generative self-supervised methods. For example, RVSA [18] leverages the vision transformer backbone pre-trained by the vanilla MAE [10] method and introduces computationally efficient rotated varied-size window attention to replace the original attention during fine-tuning. In addition, many studies are focusing on improving generative self-supervised algorithms by leveraging general image knowledge [25, 26, 28], scaling up parameter sizes [29], integrating spatio-temporal data [30, 31, 32], handling multi-sensor data [76, 77, 78, 79, 80], and employing multi-scale concepts [33, 35, 36]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
3 Methodology
The proposed pipeline contains two critical components: dataset generation and efficient pre-training, which will be introduced in detail in the following text.
3.1 Dataset Generation
@@
Figure 1: Left: The organization of data sources in OpticalRS-4M. Right: The comparison between OpticalRS-4M and other datasets.
Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets. To overcome this, we first curate a large-scale RS dataset, OpticalRS-4M, through extensive data collection and preprocessing.
3.1.1 Data Collection
We began by reviewing publicly available RS datasets from the past decade. Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [20]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity. Intra-class Diversity: High diversity within classes ensures comprehensive representation of real-world target distributions. To ensure this, we consider the data from different sources of the same scene or target. Moreover, in the RS field, the attributes of objects may vary with geographic location and imaging time, so the change detection data is also incorporated. Inter-class Diversity: To enable self-supervised pre-training to distinguish between categories effectively, we incorporated as many fine-grained categories as possible, especially those with high semantic overlap. This strategy helps the model learn robust feature representations. 2) Richness: Beyond diversity, richness is also essential. We ensure diverse content features and large sample sizes by collecting images under various conditions — weather, seasons, lighting, and so on — introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion. Our data collection process actively includes the sample of the same category under different imaging conditions to ensure comprehensive real-world representation, enhancing the model’s representation and generalization capabilities. 3) Scalability: Given the evolving applications of remote sensing images, scalability is also crucial. We constantly enrich the dataset based on scene and target categories across classification, detection, and segmentation tasks.
In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios. Fig. 1 left illustrates the included RS datasets collected.
3.1.2 Data Preprocessing
Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained. To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
1) We focused exclusively on optical images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training. 2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices. 3) We filtered out images with fewer than 64 or more than 1024 pixels. 4) The remaining images were combined and duplicates were removed using a two-phase approach: first, a coarse phase with perceptual hashing [81], and then a refined phase involving manual review. This ensured that only highly similar images were excluded. As a result of these steps, we created the OpticalRS-4M dataset, which contains approximately 4 million high-quality RS images — significantly larger than previous representative RS datasets, as illustrated in Fig. 1 right.
In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality results. We’ve screened and processed most small and medium-scale datasets from the past decade in this field. More efforts can be made to extend OpticalRS-4M by including larger datasets like SSL4EO-S12 [54] and SSL4EO-L [55] to enhance dataset capacity.
3.1.3 Dataset Diversity
Refer to captionRefer to caption
Figure 2: Left: Main categories with visual examples from OpticalRS-4M.Right: Full hierarchy of OpticalRS-4M categories.
To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, reviewing all the scenarios and targets included in OpticalRS-4M, as shown in Fig. 2. The analysis reveals that OpticalRS-4M comprises 12 main categories, each containing numerous subclasses. Unlike existing self-supervised optical remote sensing datasets, OpticalRS-4M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category. This category includes labels such as “Concert”, “Parade/Protest”, “Religious Activity”, “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and “Traffic Collision”. This expanded information enhances the versatility of OpticalRS-4M for a wider range of downstream tasks.
3.2 Efficient Pre-training
After obtaining a large-scale RS dataset, the next step is to conduct the self-supervised pre-training. However, as the capacity of the dataset grows, vast computational and time costs are required for existing MIM approaches used in the RS community, e.g., MAE [10]. To address this issue, our pipeline introduces an efficient MIM method, i.e., SelectiveMAE. In this section, we first review the preliminaries of MAE, then we present the details of SelectiveMAE.
3.2.1 Masked Autoencoders Preliminaries